# Binary VPN Traffic Detection Using Wavelet Features and Machine Learning

### v2

@@ I Introduction
The widespread adoption of network encryption poses a challenge for traffic classification, as traditional deep packet inspection and port-based methods are no longer effective [CiscoVNI2017, Statista2020]. VPNs are among the most commonly used encryption tools, with nearly one-third of internet users employing VPN services to protect their online activity [GlobalWebIndex2020]. While this enhances privacy and security, it also complicates traffic analysis for network management, intrusion detection, and policy enforcement.
Recent research has explored machine learning (ML) for encrypted traffic classification. The approaches in this field can be broadly categorized into three main groups. First, traditional feature-based methods have shown promise, with techniques like Sample Entropy Fingerprint and TCP characteristics heuristics achieving high accuracy in VPN detection [9353766, hanlon2024detecting]. Second, deep learning approaches have emerged as powerful tools, with innovations like FlowPics for visual representation of flow features [10724208] and advanced architectures combining CNN with attention mechanisms [Chai2024, Seydali2024]. The third category explores novel modeling paradigms, including multi-task learning with DistilBERT [Park2024] and text-to-text transformation using large language models [Luo2024].
[…]
@@ II-B Flow Processing and Analysis
We utilized NFStream [Aouini2022], a flexible network data analysis framework, to process the dataset’s PCAP files. While VNAT provides both traffic traces in PCAP format and pre-extracted statistical summaries in HDF (h5) format, we chose NFStream for its ability to seamlessly integrate plugins, particularly for calculating the Wavelet Features discussed in Section II-C.
@@ II-C Wavelet Features
Wavelet transform is a time-frequency analysis method that decomposes signals into components at different scales, making it suitable for analyzing non-stationary traffic patterns. Unlike Fourier transforms, wavelets preserve both time and frequency information, capturing traffic bursts and periodic behaviors effectively [Ahad2016].
@@
    Shannon Entropy ($H_{j}$) quantifies the information content [Shannon1948]:
@@ II-D Feature Extraction
For our classification task, we extract wavelet-based features to capture the temporal characteristics of packet size distributions in network flows. We apply DWT to transport-layer packet size sequences in both forward and backward directions. Following [10044382], we remove flows with fewer than 20 packets. Our choice of decomposition level J=5 is supported by the packet count distribution in our filtered dataset, as shown in Table II. With the 25th percentile at 56 packets, J=5 (scale of $322^{5}=$ 32 according to the optimal decomposition level formula $J=\lfloor\log_{2}(n)\rfloor$) ensures effective capture of packet variations across the majority of flows. We also evaluate J=12 to enable direct comparison with [10044382], though at this level (scale of $40962^{12}=$ 4096), the decomposition scale significantly exceeds typical flow lengths in their dataset. The highly skewed distribution (mean $=2,599$, median $=63$) indicates that while some flows contain up to 857,900 packets, such extensive flows are rare outliers. This distribution suggests that while J=12 might capture finer-grained variations in exceptionally long flows, it might provide diminishing returns for typical traffic patterns.
@@ IV Conclusion
This work utilizes the VNAT dataset [10044382] as the primary data source. All processed datasets derived from these traces—including extracted network flows, filtered datasets, and wavelet-based feature representations—along with analysis scripts, trained models, and performance metrics necessary to reproduce our results are available at [github-repo], ensuring full reproducibility of the machine learning experiments.

### v3

@@ I Introduction
The widespread adoption of network encryption poses a challenge for traffic classification, as traditional deep packet inspection and port-based methods are no longer effective [1, 2]. VPNs are among the most commonly used encryption tools, with nearly one-third of internet users employing VPN services to protect their online activity [3]. While this enhances privacy and security, it also complicates traffic analysis for network management, intrusion detection, and policy enforcement.
Recent research has explored machine learning (ML) for encrypted traffic classification. The approaches in this field can be broadly categorized into three main groups. First, traditional feature-based methods have shown promise, with techniques like Sample Entropy Fingerprint and TCP characteristics heuristics achieving high accuracy in VPN detection [4, 5]. Second, deep learning approaches have emerged as powerful tools, with innovations like FlowPics for visual representation of flow features [6] and advanced architectures combining CNN with attention mechanisms [7, 8]. The third category explores novel modeling paradigms, including multi-task learning with DistilBERT [9] and text-to-text transformation using large language models [10].
Notably, [11][11] introduced an ML framework that combines wavelet-based features with uncertainty quantification for application-level classification across VPN and non-VPN traffic. However, the related but distinct problem of binary VPN detection—determining whether traffic is VPN-encrypted regardless of application—has received less attention and presents unique value for network policy enforcement and security monitoring. While wavelet transforms have proven effective in network anomaly detection[12, 13], their potential for binary VPN classification remains unexplored. Building on VNAT’s approach, our study applies wavelet-based feature extraction specifically for binary VPN detection. While both studies use wavelet decomposition, our work differs from VNAT in three key aspects: objectives (application classification vs. binary VPN detection), methodology (we systematically evaluate decomposition depths J=12 and J=5, finding J=12 superior despite diminishing returns), and architecture (demonstrating that RF achieves comparable performance to VNAT’s specialized Prototypical Network).
@@ II-B Flow Processing and Analysis
We utilized NFStream [14], a flexible network data analysis framework, to process the dataset’s PCAP files. While VNAT provides both traffic traces in PCAP format and pre-extracted statistical summaries in HDF (h5) format, we chose NFStream for its ability to seamlessly integrate plugins, particularly for calculating the Wavelet Features discussed in Section II-C.
@@ II-C Wavelet Features
Wavelet transform is a time-frequency analysis method that decomposes signals into components at different scales, making it suitable for analyzing non-stationary traffic patterns. Unlike Fourier transforms, wavelets preserve both time and frequency information, capturing traffic bursts and periodic behaviors effectively [15].
@@
    Shannon Entropy ($H_{j}$) quantifies the information content [19]:
@@ II-D Feature Extraction
For our classification task, we extract wavelet-based features to capture the temporal characteristics of packet size distributions in network flows. We apply DWT to transport-layer packet size sequences in both forward and backward directions. Following [11], we remove flows with fewer than 20 packets. Our choice of decomposition level J=5 is supported by the packet count distribution in our filtered dataset, as shown in Table II. With the 25th percentile at 56 packets, J=5 (scale of $2^{5}=32$ according to the optimal decomposition level formula $J=\lfloor\log_{2}(n)\rfloor$) ensures effective capture of packet variations across the majority of flows. We also evaluate J=12 to enable direct comparison with [11], though at this level (scale of $2^{12}=4096$), the decomposition scale significantly exceeds typical flow lengths in their dataset. The highly skewed distribution (mean =2,599=2,599, median =63=63) indicates that while some flows contain up to 857,900 packets, such extensive flows are rare outliers. This distribution suggests that while J=12 might capture finer-grained variations in exceptionally long flows, it might provide diminishing returns for typical traffic patterns.
@@ IV Conclusion
This work utilizes the VNAT dataset [11] as the primary data source. All processed datasets derived from these traces—including extracted network flows, filtered datasets, and wavelet-based feature representations—along with analysis scripts, trained models, and performance metrics necessary to reproduce our results are available at [20], ensuring full reproducibility of the machine learning experiments.